# Modeling wildfire dynamics through a physics-based approach incorporating fuel moisture and landscape heterogeneity

### v1

@@ 1 Introduction
Predicting the impact of wildfires on our ecosystems is challenging because wildfire propagation depends on multiple factors and processes across multiple scales, each addressing different physical and chemical processes involved in the combustion and fire propagation phenomena. The knowledge of these scales is crucial for the formulation of a wildfire propagation model. Depending on the purpose of the model, certain scales are more suitable than others. If we seek a model with predictive capacity—i.e., one that operates faster than real time [23, 27, 28, 50, 52, 54, 61]—it is not feasible for the model to resolve all scales. Instead, only the larger scales are explicitly represented in the model, while the processes occurring at the smaller scales must be approximated or modeled indirectly. Resolving these small-scale phenomena would require a very fine spatial discretization, resulting in prohibitively high computational costs.
Following the criteria and nomenclature proposed by Sero et al. [30, 51], we consider four relevant scales: the gigascopic scale, the macroscopic scale, the mesoscopic scale and the microscopic scale. The gigascopic scale encompasses fire spread over large areas, several hundred meters to kilometers. At this scale, the fire front can be represented as a one-dimensional curve advancing across the landscape. If we now look more closely at the vegetation stratum, we can define the macroscopic scale modeling the vegetation layer, the air above and the ground underneath. We define an occupation density or surface coverage (SC) and determine the width and height of the fire front. A one level finer scale is the mesoscopic scale, which allows us to characterize the fine structure of vegetation, such as the geometry of branches and leaves. At this scale, the vegetation stratum is regarded as a two-phase porous medium, consisting of a mixture of vegetation (solid) and air (gas). Finally, the smallest scale is the microscopic scale, which allows us to define individual components of vegetation, including solid, liquid, and gaseous phases. At this scale, we can observe pyrolysis and drying, describe how wood is converted into by-products such as char and flammable gases.
@@
In this work, we extend the ADR wildfire propagation model presented in [42] to explicitly represent fuel moisture effects, distinguishing between live and dead fuel moisture content when desired. We propose a simple approach, based on the apparent calorific capacity method, which models the effect of thermal phase changes by considering an apparent (or effective) heat capacity, usually defined as a piecewise constant function of the temperature [14, 55]. Three different moisture models, representing the evaporation process and the constituents of fuel with a different level of detail, are presented. For a more complete model, we represent the underlying physio-chemical mechanisms involved in live and dead fuel moisture balance to provide an approximation of the specific heat for the live and dead fuel independently.
@@
The paper is structured as follows: In Sec. 2, the variables of the model are defined, and the differential equations are derived from physical balance laws, with additional details provided in App. A. We discuss modeling decisions on the wind, radiation, combustion, and, introduce different moisture models in Sec. 2.6. We present the numerical schemes in Sec. 3, accompanied by validation results in App. B. Sec. 4 investigates the effects of the discussed mechanisms by numerical simulations. Two-dimensional simulations in Sec. 4.5 lead the path to more realistic scenarios. Finally, Sec. 5 summarizes the highlights, critically discusses short-comings and provides an outlook on future challenges.
@@ 2 Mathematical description of the model
After discussing the representation of the vegetation layer, we derive the advection diffusion reaction model by physical balance laws. Variants of different terms included follow, and we close this section with fixing physical parameters for the simulations. In App. A, a detailed derivation of the equations for the conservation of energy, which complements this section, can be found.
@@ 2.1 Representation of the vegetation layer
We consider a wildfire propagation model, defined in the spatial domain $\Omega=\Lambda\times[0,l]$, with $\Lambda\subset\mathbb{R}^{2}$ as the ground surface area and $l$ as the height. Let $V=S\times[0,l]\subset\Omega$ be a control volume inside this domain, with $S\subset\Lambda$ being the ground surface area of this control volume. Within $V$, we take a continuum approach to describe the biomass and air mixture and use bulk parameters to characterize the properties of this continuum. We assume that all modeled quantities are two-dimensional—homogenized in $l$ over $S$. Thus, all variables of interest will be defined on $S$, with $\mathbf{x}=(x,y)\in\Lambda$ denoting the spatial location. Further, within $V$, we assume a mixture of two phases: (i) the solid phase, composed of solid fuel (that we will consider equal to the biomass), and (ii) the gaseous phase, which consists of air and combustion products [30, 51, 52]. The mixture can be modeled as a two-phase porous medium [10, 59]. We then define the following bulk parameters for quantities within $V$ at the centroid $\mathbf{x}_{c}$ of $S$: solid fuel with volume $V_{f}$, density $\bar{\rho}_{f}$, the specific heat $\bar{c}_{p,f}$ and (dry) mass $m_{f}=\bar{\rho}_{f}V_{f}$, as well as air and other combustion products with volume $V_{a}$, density $\bar{\rho}_{a}$, specific heat $\bar{c}_{p,a}$ and mass $m_{a}=\bar{\rho}_{a}V_{a}$, that is $m=m_{f}+m_{a}$. We assume that the fuel density and specific heat may depend upon space, that is to say $\bar{\rho}_{f}=\bar{\rho}_{f}(\mathbf{x})$ and $\bar{c}_{p,f}=\bar{c}_{p,f}(\mathbf{x})$, whereas the density and specific heat of air and combustion products, $\bar{\rho}_{a}$ and $\bar{c}_{p,a}$, are spatially constant.
@@
where $W_{f,0}(\mathbf{x})=m_{f,0}(\mathbf{x})/\lvert S\rvert$ is the initial fuel load per surface unit, with $m_{f,0}$ the initial fuel mass distribution [3]. The initial fuel load $W_{f,0}$ usually depends upon the surface coverage, $SC\in[0,1]$, which can be obtained from remote sensing techniques. We express the initial fuel load as
@@
The fuel mass fraction relates to the quantities defined above as
@@
where $m_{f,0}$ and $V_{f,0}$ are the initial mass and volume of the fuel.
In what follows, we drop the dependency of the variables upon $\mathbf{x}$ and $t$ for the sake of simplicity. The expression to compute the bulk density from the fuel mass fraction is obtained from Eqs. (1) and (5) and reads [59]
@@
Analogously, we can compute the bulk specific heat inside $V$ as [59]
@@
The thickness $l$ in the vertical direction can be seen as a relative height (with unit), herein assumed to be constant $l=1$ for the sake of simplicity. In the vertically averaged system, the vertical component is not modeled directly, and in the following derivation of the model, we will highlight where the simplification from a three-dimensional to a two-dimensional model comes into play.
@@ 2.2 Physical derivation of the ADR model
Let us consider a fixed volume defined as $V=S\times[0,l]\subset\Omega$, with $S\subset\Lambda$ the ground layer, and with $l$ the thickness in the vertical direction. The boundary of the control volume is denoted by $\partial V$. In what follows, we will assume that there is a local thermal equilibrium in the medium, that is the temperature of the solid fuel, $T_{f}$, and the temperature of the gases, $T_{a}$, will coincide, $T_{f}=T_{a}=T$ [51, 52]. Besides, we consider that the effect of pressure variations and viscous dissipation can be neglected as shown in App. A. The conservation of energy inside $V$ can be expressed as (see App. A, Eq. (95))
@@
$\frac{d}{dt}\int_{V}\rho h\,\mathrm{d}V+\int_{\partial V}\bar{\rho}_{a}h_{a}(1-R_{f})\mathbf{v}_{a}\cdot\mathbf{n}\,\mathrm{d}\Gamma=\dot{{Q}}_{\mathrm{cond}}-\dot{{Q}}_{\mathrm{conv}}+\dot{{Q}}_{\mathrm{rad}}+\dot{{Q}}_{\mathrm{reac}}\,,$ (10)
@@
where $h$ is the specific enthalpy and $\rho$ is the density of the mixture, i.e. $\rho h=\bar{\rho}_{f}h_{f}R_{f}+\bar{\rho}_{a}h_{a}(1-R_{f})$, with $h_{f}$ and $h_{a}$ being the specific enthalpy of the solid and gas phase, respectively. $\dot{{Q}}_{\mathrm{cond}}$, $\dot{{Q}}_{\mathrm{conv}}$ and $\dot{{Q}}_{\mathrm{rad}}$ are the conduction, vertical convection and radiation heat fluxes, respectively, and $\dot{{Q}}_{\mathrm{reac}}$ is the reaction heat source due to combustion. These terms will be modelled following the approaches often used in the literature [4, 13, 27, 28, 42, 61]. As mentioned before, all variables are homogenized in the vertical direction, $z$, thus the dependency of the variables on $z$ is dropped and gradients in $z$-direction are assumed zero. We only consider the $x$- and $y$-component in vector variables and the spatial differential operator will be defined as $\nabla=(\partial/\partial x,\partial/\partial y)^{T}$.
@@
The advection term accounts for the advection of energy of the gaseous phase, $\bar{\rho}_{a}h_{a}$, and is governed by the so-called Darcy velocity or seepage velocity $\mathbf{v}_{D}$ in the context of porous flows [59]
$\mathbf{v}_{D}=(1-R_{f})\mathbf{v}_{a},$ (11)
@@
where $\mathbf{v}_{a}$ is the actual flow velocity through the solid fuel. We define a bulk advection velocity, $\mathbf{v}=(u,v)^{T}$, that allows to rewrite the advection term in Eq. (10) more compactly. The bulk velocity $\mathbf{v}$ is related to the actual flow velocity through the solid fuel, $\mathbf{v}_{a}$, as
@@
We rewrite Eq. (10) as
$\frac{d}{dt}\int_{V}H\,\mathrm{d}V+\int_{\partial V}H\mathbf{v}\cdot\mathbf{n}\,\mathrm{d}\Gamma=\dot{{Q}}_{\mathrm{cond}}-\dot{{Q}}_{\mathrm{conv}}+\dot{{Q}}_{\mathrm{rad}}+\dot{{Q}}_{\mathrm{reac}},$ (14)
with $H=\rho h$ the enthalpy per unit volume and $h=h(T)$ the specific enthalpy, which depends on the temperature.
@@
where $\dot{\mathbf{q}}$ is the heat flux vector and $k_{c}$ is the thermal conductivity of the bulk mixture. The convection heat flux is given by Newton’s cooling law
$\dot{{Q}}_{\mathrm{conv}}=\int_{S}\hat{\alpha}(T-T_{\infty})\,\mathrm{d}\Gamma,$ (16)
with $\hat{\alpha}$ the convection coefficient and $T_{\infty}$ the temperature of the air over the surface.
In this model, we assume that radiation is only acting locally by assuming that the medium is optically thick. Following the Rosseland approximation, the radiation heat flux can be written in the form of a nonlinear conduction heat flux as [5, 27]
@@
provided that the optical path length for radiation, $\delta$, is smaller than the characteristic lengths of the modeled physical phenomena. Other parameters in the previous expression are the Stefan-Boltzmann constant, $\sigma$, and the emissivity factor, $\epsilon$.
@@
$\dot{{Q}}_{\mathrm{reac}}=-\int_{V}\bar{\rho}_{f}R_{f,0}\dot{Y}\mathcal{H}\,\mathrm{d}V\,,$ (18)
where $\mathcal{H}$ is the combustion heat per unit mass of fuel and $\dot{Y}$ is the fuel consumption rate per unit volume, given by a first-order kinetic equation
$\dot{Y}:=\frac{\partial Y}{\partial t}=-\Psi(T)Y$ (19)
with the reaction rate $\Psi(T)$ being discussed in Sec. 2.5. By inserting Eq. (19) into Eq. (18), we obtain
$\dot{{Q}}_{\mathrm{reac}}=\int_{V}\bar{\rho}_{f}R_{f,0}\mathcal{H}\Psi(T)Y\,\mathrm{d}V.$ (20)
Due to the dimensionality reduction, advection in the vertical direction is not accounted for, thus, we can have $\nabla\cdot\mathbf{v}\neq 0$, i.e. the incompressibility condition is not guaranteed. In this case, some correction should be considered to avoid the accumulation/disappearance of energy at the sources/sinks of the advection velocity field [37]. A simple approach would be to consider an additional vertical convection flux
$\dot{{Q}}_{\mathrm{conv}}^{\prime}=\int_{S}H\hat{w}\,\mathrm{d}\Gamma,$ (21)
where $\hat{w}$ is the (modeled) wind in the vertical direction satisfying $\frac{\partial u}{\partial x}+\frac{\partial v}{\partial y}+\frac{\partial\hat{w}}{\partial z}=0$, i.e., $\hat{w}=-\frac{1}{S}\int_{\partial V}\mathbf{v}\cdot\mathbf{n}\,\mathrm{d}\Gamma$. For the sake of simplicity, this is not considered in the model presented here.
Using the Gauß divergence theorem, we rewrite Eq. (14) as
@@
where $\alpha=\hat{\alpha}/l$, $w=\hat{w}/l$ and
@@
The relation between $T$ and $H$ in Eq. (24) is given by a function $T=T(H)$ that may be multi-valued to model phase change due to the evaporation of fuel moisture [6, 23]. In the literature, this is referred to as the Stefan problem, originally introduced for the computation of phase change materials in closed systems [22, 55, 62].
Remark 1.
The wildfire model in Eq. (24) is written in a conservative form, as it is based on the conservation of enthalpy, see Eq. (10). As a modelling simplification, we assume that the system is closed and we do not consider the loss of energy due to a loss of mass (water vapor and other volatile products and gases), which is not explicitly included. After a heating-cooling process, the final enthalpy of the described system should be smaller than its initial enthalpy due to the loss of mass resulting from the evaporated water. However, Eq. (24), without additional modifications, would not reproduce this reduction of enthalpy due to the aforementioned modelling assumptions.
Remark 1 is connected to two other aspects that are relevant in the model refinement. First, the phase change (fuel moisture evaporation) only occurs in one direction, since the water vapor then leaves the domain. Therefore, in the cooling-to-ambient-temperature process, the enthalpy-temperature relation $T=T(H)$ should be modified to prevent from a physically unfeasible latent cooling (water vapor condensation), as resolved in [5]. Second, since the advection term is written in conservative form, there may be physically unfeasible enthalpy accumulation in some spatial regions caused by a non-zero divergence of the velocity field. A correction based on Eq. (21) can be used. Recall that this non-zero divergence may be due to the dimensionality reduction of the problem, as well as to the representation of the effect of the topography proposed here (see Sec. 2.3). Taking into account the degree of simplification used here and the previous observations, a definition of the problem in full conservation form, as in Eq. (24), is not required for the purpose of this paper.
We can find alternative approaches in the literature to the multi-valued enthalpy operator method in Eq. (24) to include the effect of the phase change. Most of these approaches were developed for other engineering applications. Nevertheless, some of the difficulties mentioned above are likely to remain. The so-called heat source method separates the enthalpy into two terms, one corresponding to the sensible heat and the other to the latent heat. The latter would be included as a source term in the classical heat equation [14, 52, 55]. Another approach is called apparent calorific capacity method, which models the effect of thermal phase changes by considering an apparent (or effective) heat capacity, usually defined as a piecewise constant function [14, 55]. This approach exploits the relation $\frac{\partial H}{\partial t}=\frac{\partial H}{\partial T}\frac{\partial T}{\partial t}$, allowing to write
@@
Motivated by the observations mentioned above, we propose to use a version of Eq. (24) in non-conservative form. Using the apparent calorific capacity approach, we give an alternative formulation for the energy equation (see App. A)
@@
We introduce an additional simplification by considering fuel properties at the initial time, i.e. $\rho=\rho_{0}$ and $c_{p}=c_{p0}$ according to Eq. (8). When using the apparent capacity approach, the complete model reads
@@
Considering $R_{f,0}=1$, $\bar{c}_{p,f}=\bar{c}_{p,a}$ and $\bar{\rho}_{f}=\bar{\rho}_{a}$ as constant values, we recover the model in [42].

### v2

@@ 1 Introduction
Predicting the impact of wildfires on our ecosystems is challenging because wildfire propagation depends on multiple factors and processes across multiple scales, each addressing different physical and chemical processes involved in the combustion and fire propagation phenomena. Knowledge of these scales is crucial for the formulation of a wildfire propagation model. Depending on the purpose of the model, certain scales are more suitable than others. When seeking a model with predictive capacity—i.e. one that operates faster than real time [25, 29, 30, 53, 55, 57, 64]—it is not feasible for the model to resolve all scales. Instead, only the larger scales are explicitly represented in the model, while the processes occurring at the smaller scales must be approximated or modeled indirectly. Resolving these small-scale phenomena would require a very fine spatial discretization, resulting in prohibitively high computational costs.
Following the criteria and nomenclature proposed by Sero et al. [32, 54], there are four relevant scales: the gigascopic scale, the macroscopic scale, the mesoscopic scale and the microscopic scale. The gigascopic scale encompasses fire spread over large areas, several hundred meters to kilometers. At this scale, the fire front can be represented as a one-dimensional curve advancing across the landscape. Looking more closely at the vegetation stratum, the macroscopic scale is defined to model the vegetation layer, the air above and the ground underneath. On this scale, the occupation density or surface coverage (SC) is defined and the width and height of the fire front are determined. A one-level finer scale is the mesoscopic scale, which allows us to characterize the fine structure of vegetation, such as the geometry of branches and leaves. At this scale, the vegetation stratum is regarded as a two-phase porous medium, consisting of a mixture of vegetation (solid) and air (gas). Finally, the smallest scale is the microscopic scale, which enables the definition of individual components of vegetation, including solid, liquid, and gaseous phases. At this scale, pyrolysis and drying can be represented and the process by which wood is converted into by-products such as char and flammable gases is described.
@@
In this work, we extend the ADR wildfire propagation model presented in [45] to explicitly represent fuel moisture effects, distinguishing between live and dead fuel moisture content when desired. The proposed approach is based on the apparent calorific capacity method, which models the effect of thermal phase changes by considering an apparent (or effective) heat capacity, usually defined as a piecewise constant function of the temperature [14, 58]. Three different moisture models, representing the evaporation process and the constituents of fuel with a different level of detail, are presented. The most comprehensive model captures the underlying physio-chemical mechanisms involved in the moisture balance of both live and dead fuel, allowing for an approximation of their specific heat independently.
@@
The paper is structured as follows: In Sec. 2, the variables of the model are defined, and the differential equations are derived from physical balance laws, with additional details provided in App. A. We discuss modeling decisions on the wind, radiation, combustion, and, introduce different moisture models in Sec. 2.6. The numerical schemes are presented in Sec. 3, accompanied by verification results in App. B. Sec. 4 investigates the effects of the discussed mechanisms by numerical simulations. Two-dimensional simulations in Sec. 4.5 lead the path to more realistic scenarios. Finally, Sec. 5 summarizes the highlights, critically discusses shortcomings and provides an outlook on future challenges.
@@ 2 Mathematical description of the model
After discussing the representation of the vegetation layer, the advection diffusion reaction model is derived by physical balance laws. Variants of different terms included follow, and this section closes with fixing physical parameters for the simulations. In App. A, a detailed derivation of the equations for the conservation of energy, which complements this section, can be found.
@@ 2.1 Representation of the vegetation layer
The wildfire propagation model is defined in the spatial domain $\Omega=\Lambda\times[0,l]$, with $\Lambda\subset\mathbb{R}^{2}$ as the ground surface and $l$ [m] as the height. Let $V=S\times[0,l]\subset\Omega$ be a control volume inside this domain, with $S\subset\Lambda$ being the ground layer of this control volume. Within $V$, a continuum approach is used to describe the biomass and air mixture, allowing to define bulk parameters to characterize the properties of this continuum. Averaging all modeled quantities in $l$ over $S$ leads to two dimensional space-dependencies. Thus, all variables of interest will be defined on $S$, with $\mathbf{x}=(x,y)\in\Lambda$ denoting the spatial location. A mixture of two phases is assumed within $V$: (i) the solid phase, composed of solid fuel (that is considered equal to the biomass), and (ii) the gaseous phase, which consists of air and combustion products [32, 54, 55]. The mixture can be modeled as a two-phase porous medium [10, 62]. The abstraction from the two-phase porous medium leads to the definition of bulk parameters for quantities within $V$ at the centroid $\mathbf{x}_{c}$ of $S$: solid fuel with volume $V_{f}$ [m3], density $\bar{\rho}_{f}$ [kg/m3], the specific heat $\bar{c}_{p,f}$ [kJ/(kg·K)] and (dry) mass $m_{f}=\bar{\rho}_{f}V_{f}$ [kg], as well as air and other combustion products with volume $V_{a}$ [m3], density $\bar{\rho}_{a}$ [kg/m3], specific heat $\bar{c}_{p,a}$ [kJ/(kg·K)] and mass $m_{a}=\bar{\rho}_{a}V_{a}$ [kg], that is $m=m_{f}+m_{a}$. The fuel density and specific heat may depend upon space, that is to say $\bar{\rho}_{f}=\bar{\rho}_{f}(\mathbf{x})$ and $\bar{c}_{p,f}=\bar{c}_{p,f}(\mathbf{x})$, whereas the density and specific heat of air and combustion products, $\bar{\rho}_{a}$ and $\bar{c}_{p,a}$, are assumed to be spatially constant.
@@
where $W_{f,0}(\mathbf{x})=m_{f,0}(\mathbf{x})/\lvert S\rvert$ is the initial fuel load per surface unit, with $m_{f,0}$ the initial fuel mass distribution [3]. The initial fuel load $W_{f,0}$ [kg/m2] usually depends upon the surface coverage, $SC\in[0,1]$, which can be obtained from remote sensing techniques. The initial fuel load is expressed as
@@
The remaining fuel mass fraction relates to the quantities defined above as
@@
where $m_{f,0}$ and $V_{f,0}$ are the initial mass and volume of the fuel. Note that $Y$ will hereafter be referred to as the remaining fuel fraction or the remaining biomass fraction for the sake of brevity.
In what follows, the dependency of the variables upon $\mathbf{x}$ and $t$ is dropped for the sake of simplicity. The expression to compute the bulk density from the remaining fuel fraction is obtained from Eqs. (1) and (5) and reads [62]
@@
Analogously, the bulk specific heat inside $V$ is computed as follows [62]
@@
The thickness $l$ in the vertical direction can be seen as a relative height (with unit), herein assumed to be constant $l=1$ m for simplicity. In the vertically averaged system, the vertical component is not modeled directly, and in the following derivation of the model, the simplification from a three-dimensional to a two-dimensional model is highlighted.
@@ 2.2 Physical derivation of the ADR model
Consider a fixed control volume defined as $V=S\times[0,l]\subset\Omega$, with $S\subset\Lambda$ the ground layer, and with $l$ the thickness in the vertical direction. The control volume is open along the lateral and top boundaries. In what follows, it is assumed that the fuel consumption rate is given by the following phenomenological equation [5, 54]
$\dot{Y}:=\frac{\partial Y}{\partial t}=-\Psi(T)Y$ (10)
which represents a first-order kinetic equation with the reaction rate $\Psi(T)$ being discussed in Sec. 2.5. A highly simplified chemical reaction is considered, where the fuel is assumed to be the only reactive specie, and the product is a gas with properties similar to air. It is assumed that all the fuel undergoes complete combustion and is fully converted into gas. Thus, the product production rate $\dot{m}_{a}^{R}$ [kg/s] is equal to the fuel consumption rate $\dot{m}_{f}={m}_{f,0}\dot{Y}$ [kg/s], so $\dot{m}_{a}^{R}=-\dot{m}_{f}$.
To derive the ADR model, we apply the equations for the conservation of mass and energy within the selected control volume. As mentioned before, all variables are homogenized in the vertical direction, $z$, thus the dependency of the variables on $z$ is dropped and gradients in $z$-direction are assumed zero. All vector variables only have the $x$- and $y$-component and the spatial differential operator is defined as $\nabla=(\partial/\partial x,\partial/\partial y)^{T}$. In addition, incompressibility conditions are assumed within the mixture. The equation for the conservation of mass inside $V$ can be expressed as
$\frac{d}{dt}\int_{V}\rho\,\mathrm{d}V+\int_{\partial V}\bar{\rho}_{a}(1-R_{f})\mathbf{v}_{a}\cdot\mathbf{n}\,\mathrm{d}\Gamma=-\dot{m}_{\mathrm{out}}\,,$ (11)
where $\rho=\bar{\rho}_{f}R_{f}+\bar{\rho}_{a}(1-R_{f})$ is the density of the mixture, $\mathbf{v}_{a}=(u_{a},v_{a})^{T}$ [m/s] is the actual (2D) velocity of the gas phase through the solid fuel and $\dot{m}_{\mathrm{out}}$ is the mass flow leaving the domain through the top layer, exiting into the atmosphere. Note that $\partial V$ represents the boundary surfaces of the control volume, excluding the top and bottom layers. Using the relation $\dot{R_{f}}=\dot{Y}R_{f,0}$ on the first term of Eq. (11), we obtain
$\dot{m}_{\mathrm{out}}=\dot{Y}R_{f,0}(\bar{\rho}_{a}-\bar{\rho}_{f})V-\int_{\partial V}\bar{\rho}_{a}(1-R_{f})\mathbf{v}_{a}\cdot\mathbf{n}\,\mathrm{d}\Gamma.$ (12)
Note that the mass flow leaving the domain is due to two contributions: (i) the first term on the right-hand side of Eq. ((12)) represents the excess mass of gas that cannot occupy the volume vacated by the fuel upon its consumption, it will be zero without combustion; (ii) the second term on the right-hand side of Eq. (12) represents the mass flow difference through the boundaries of the domain in 2D, due to variations in $R_{f}$ and a divergence of $\mathbf{v}_{a}$. Assuming that $R_{f}\lesssim 0.01$ [64], the approximation $1-R_{f}\approx 1$ allows to rewrite Eq. (12) as
$\dot{m}_{\mathrm{out}}\approx\dot{Y}R_{f,0}(\bar{\rho}_{a}-\bar{\rho}_{f})V-\bar{\rho}_{a}\int_{\partial V}\mathbf{v}_{a}\cdot\mathbf{n}\,\mathrm{d}\Gamma.$ (13)
Note that the second term will be zero when the incompressibility condition is satisfied in the 2D advection field, i.e. $\nabla\cdot\mathbf{v}_{a}=0$, leading to
$\dot{m}_{\mathrm{out}}=\dot{Y}R_{f,0}(\bar{\rho}_{a}-\bar{\rho}_{f})V$ (14)
@@
Now consider the equation for the conservation of energy. Assume that there is a local thermal equilibrium in the medium, that is, the temperature of the solid fuel, $T_{f}$ [K], and the temperature of the gases, $T_{a}$ [K], will coincide, $T_{f}=T_{a}=T$ [54, 55]. In addition, the effect of pressure variations and viscous dissipation is neglected, as shown in App. A. The conservation of energy inside $V$ can be expressed as (see App. A, Eq. (100))
@@
$\frac{d}{dt}\int_{V}\rho h\,\mathrm{d}V+\int_{\partial V}\bar{\rho}_{a}h_{a}(1-R_{f})\mathbf{v}_{a}\cdot\mathbf{n}\,\mathrm{d}\Gamma=\dot{{Q}}_{\mathrm{cond}}+\dot{{Q}}_{\mathrm{rad}}-\dot{{Q}}_{\mathrm{out}}+\dot{{Q}}_{\mathrm{comb}}\,,$ (15)
@@
where $h$ [kJ/kg] is the specific enthalpy and $\rho$ [kg/m3] is the density of the mixture, i.e. $\rho h=\bar{\rho}_{f}h_{f}R_{f}+\bar{\rho}_{a}h_{a}(1-R_{f})$, with $h_{f}$ and $h_{a}$ being the specific enthalpy of the solid and gas phase, respectively. $\dot{{Q}}_{\mathrm{cond}}$, $\dot{{Q}}_{\mathrm{rad}}$ and $\dot{{Q}}_{\mathrm{out}}$ [kJ/s] are the conduction, radiation and free cooling (atmospheric exchange) heat fluxes, respectively, and $\dot{{Q}}_{\mathrm{comb}}$ [kJ/s] is the reaction heat source due to combustion. These terms will be modelled following the approaches often used in the literature [4, 13, 29, 30, 45, 64].
@@
The advection term accounts for the advection of energy of the gaseous phase, $\bar{\rho}_{a}h_{a}$, and can be rewritten in terms of the so-called Darcy velocity $\mathbf{v}_{D}$ in the context of porous flows [62]
$\mathbf{v}_{D}=(1-R_{f})\mathbf{v}_{a}.$ (16)
@@
The enthalpy-weighed bulk advection velocity, $\mathbf{v}=(u,v)^{T}$, allows to rewrite the advection term in Eq. (15) more compactly. The bulk velocity $\mathbf{v}$ is related to the actual flow velocity through the solid fuel, $\mathbf{v}_{a}$, as
@@
Rewriting Eq. (15) with $H=\rho h$ the enthalpy per unit volume and $h=h(T)$ the specific enthalpy, which depends on the temperature, gives
$\frac{d}{dt}\int_{V}H\,\mathrm{d}V+\int_{\partial V}H\mathbf{v}\cdot\mathbf{n}\,\mathrm{d}\Gamma=\dot{{Q}}_{\mathrm{cond}}+\dot{{Q}}_{\mathrm{rad}}-\dot{{Q}}_{\mathrm{out}}+\dot{{Q}}_{\mathrm{comb}}.$ (19)
@@
where $\dot{\mathbf{q}}$ [kJ/(m2·s)] is the heat flux vector and $k_{c}$ [kJ/(m·K·s)] is the thermal conductivity of the bulk mixture.
In this model, radiation is assumed to act only locally by assuming that the medium is optically thick. Following the Rosseland approximation, the radiation heat flux can be written in the form of a non-linear conduction heat flux as [5, 29]
@@
provided that the optical path length for radiation, $\delta$ [m], is smaller than the characteristic lengths of the modeled physical phenomena. Other parameters in the previous expression are the Stefan-Boltzmann constant, $\sigma$ [kJ/(m2·K4·s)], and the emissivity factor, $\epsilon$.
The free cooling heat flux can be expressed as
$\dot{{Q}}_{\mathrm{out}}=\dot{{m}}_{\mathrm{out}}h_{a}+\dot{{Q}}_{\mathrm{out}}^{\prime},$ (22)
where $\dot{{m}}_{\mathrm{out}}h_{a}=\dot{{m}}_{\mathrm{out}}\bar{c}_{p,a}T$ represents the energy flux into the atmosphere due to the escape of mass from the control volume and $\dot{{Q}}_{\mathrm{out}}^{\prime}$ models other heat transfer phenomena between the system and the atmosphere (natural convection, radiation, etc.). The term $\dot{{m}}_{\mathrm{out}}h_{a}$ can be expressed as a flux of the bulk enthalpy
$\dot{{m}}_{\mathrm{out}}h_{a}=H\left(\dot{Y}R_{f,0}V\left(\frac{\bar{\rho}_{a}-\bar{\rho}_{f}}{\rho}\right)-\int_{\partial V}\frac{\bar{c}_{p}}{\bar{c}_{p,a}}\mathbf{v}\cdot\mathbf{n}\,\mathrm{d}\Gamma\right).$ (23)
This term can be regarded as a generalization of the approach proposed in [40]. For the sake of simplicity, let us assume
$\dot{{Q}}_{\mathrm{out}}=\int_{S}\hat{\alpha}(T-T_{\infty})\,\mathrm{d}\Gamma,$ (24)
with $\hat{\alpha}$ [kJ/(m2·K·s)] a global heat transfer coefficient coefficient, retained as a calibration parameter, and $T_{\infty}$ [K] the temperature of the air far above the control volume.
@@
$\dot{{Q}}_{\mathrm{comb}}=-\int_{V}\bar{\rho}_{f}R_{f,0}\dot{Y}\mathcal{H}\,\mathrm{d}V\,,$ (25)
where $\mathcal{H}$ [kJ/kg] is the combustion heat per unit mass of fuel and $\dot{Y}$ is the fuel consumption rate per unit volume, given by Eq. (10). Inserting Eq. (10) into Eq. (25) gives
$\dot{{Q}}_{\mathrm{comb}}=\int_{V}\bar{\rho}_{f}R_{f,0}\mathcal{H}\Psi(T)Y\,\mathrm{d}V.$ (26)
Using the Gauß divergence theorem, Eq. (19) can be rewritten as
@@
where $\alpha=\hat{\alpha}/l$ and
@@
The relation between $T$ and $H$ in Eq. (29) is given by a function $T=T(H)$ that may be multi-valued to model phase change due to the evaporation of fuel moisture [6, 25]. In the literature, this is referred to as the Stefan problem, originally introduced for the computation of phase change materials in closed systems [24, 58, 65]. Two aspects must be taken into account in this regard. First, the phase change (fuel moisture evaporation) only occurs in one direction, since the water vapor then leaves the domain. Therefore, in the cooling-to-ambient-temperature process, the enthalpy-temperature relation $T=T(H)$ should be modified to prevent physically unfeasible latent cooling (water vapor condensation), as resolved in [5]. Second, since the advection term is written in conservative form, there may be physically unfeasible enthalpy accumulation in some spatial regions caused by a non-zero divergence of the velocity field. A correction based on Eq. (22) can be used. Recall that this non-zero divergence may be due to the dimensionality reduction of the problem, as well as to the representation of the effect of the topography proposed here (see Sec. 2.3). Taking into account the degree of simplification used here and the previous observations, a definition of the problem in full conservation form, as in Eq. (29), is not required for the purpose of this paper.
We can find alternative approaches in the literature to the multi-valued enthalpy operator method in Eq. (29) to include the effect of the phase change. Most of these approaches were developed for other engineering applications. However, some of the difficulties mentioned above are likely to remain. The so-called heat source method separates the enthalpy into two terms, one corresponding to the sensible heat and the other to the latent heat. The latter would be included as a source term in the classical heat equation [14, 55, 58]. Another approach is called apparent calorific capacity method, which models the effect of thermal phase changes by considering an apparent (or effective) heat capacity, usually defined as a piecewise constant function [14, 58]. This approach exploits the relation $\frac{\partial H}{\partial t}=\frac{\partial H}{\partial T}\frac{\partial T}{\partial t}$, allowing to write
@@
Motivated by the observations mentioned above, we propose to use a version of Eq. (29) in non-conservative form. The apparent calorific capacity approach gives an alternative formulation for the energy equation (see App. A)
@@
Considering fuel properties at the initial time, i.e. $\rho=\rho_{0}$ and $c_{p}=c_{p0}$ according to Eq. (8), simplifies the equation further. When using the apparent capacity approach, the complete model reads
@@
Considering $R_{f,0}=1$, $\bar{c}_{p,f}=\bar{c}_{p,a}$ and $\bar{\rho}_{f}=\bar{\rho}_{a}$ as constant values, the model in [45] is recovered.